OpenStudy (anonymous):

how do i find range of values of x for which 100e^-x<5

OpenStudy (unklerhaukus):

\[100e^{-x}<5\\ e^{-x}<1/20\\ -x<\ln(1/20)\\ x>-\ln(1/20)\\ x>\ln(20)\]

jhonyy9 (jhonyy9):

do you understand it why ?

OpenStudy (anonymous):

yes thank you I understand. Because the question said range I thought there should be more than one answer for x. thanks